# ACT: Agentic Classification Tree

### v2

@@ Abstract
When used in high-stakes settings, AI systems are expected to produce decisions that are transparent, interpretable, and auditable—a requirement increasingly expected by regulations. Decision trees such as CART provide clear and verifiable rules, but they are restricted to structured tabular data and cannot operate directly on unstructured inputs such as text. In practice, large language models (LLMs) are widely used for such data, yet prompting strategies such as chain-of-thought or prompt optimization still rely on free-form reasoning, limiting their ability to ensure trustworthy behaviors. We present the Agentic Classification Tree (ACT), which extends decision-tree methodology to unstructured inputs by formulating each split as a natural-language question, refined through impurity-based evaluation and LLM feedback via TextGrad. Experiments on text benchmarks show that ACT matches or surpasses prompting-based baselines while producing transparent and interpretable decision paths.
Unstructured Data: Tuberculosis diagnosis based on symptoms descriptions Patient 1: ”My main symptom is a runny nose with increased sweating, pain, fever, skin rashes, nasal congestion, sore throat, muscle pain, loss of appetite, and chills. The heavy pain is in my left temple, and I have pink rashes on the back of my neck.” Patient 2: ”I am coughing blood with pain (heavy in my left temple), skin lesions, nasal congestion, extreme fatigue affecting activities, diffuse muscle pain, loss of appetite, chills, a heavy pain on my left temple, and pink rash on the right side of my neck.” Refer to caption Refer to caption
[…]
Does this example involve coughing
up blood or weight loss?
Does the example show coughing up blood
with fever or intense fatigue?
TB
Does the example show evidence of severe
pain and either shortness of breath
or swelling in unusual locations?
TB
Not TB
Does the example show fever
and swollen lymph nodes?
Does the example show itchy or
runny nose or itchy eyes?
Not TB
TB
Not TB
yesyesnoyesnonoyesyesnono
Figure 1: Example ACT decision tree for tuberculosis diagnosis using unstructured, free-text patient descriptions. A tree is automatically learned, with each node containing a binary natural language question, autonomously discovered via recursive prompt refinement to maximize label separation at each split. At inference, these questions are answered by a large language model (LLM) from the root node to the leaves of the tree. The final classification (TB or Not TB) corresponds to the majority label of training examples described by each leaf.
@@ 1 Introduction
As AI systems become increasingly integrated into high-stakes domains such as healthcare, education, legal decision-making, and finance, the need for transparency, interpretability, and auditability in AI decision-making has intensified. These requirements are grounded not only in practical considerations, but also in legal obligations: recent governance frameworks and regulations such as the EU AI Act [^1] https://artificialintelligenceact.eu/, the OECD AI Principles [^2] https://oecd.ai/en/ai-principles, and the NIST AI Risk Management Framework [^3] https://www.nist.gov/itl/ai-risk-management-framework emphasize that AI-based decisions in high-stakes scenarios must be explainable and subject to human oversight. In this context, there is a growing need for models whose decision processes can be inspected, verified, and understood—not only by technical experts but also by stakeholders, auditors, and regulators.
@@
Historically, complex decision-making tasks in high-stakes domains—such as medical diagnosis, fraud detection, or credit underwriting—have been addressed on interpretable AI systems such as expert systems (Jackson, 1986; Shortliffe, 1986; Leonard, 1993; Talebzadeh et al., 1995) and decision trees (Breiman et al., 1984). These models allow users to trace each decision through a sequence of explicit, human-understandable rules, directly supporting the requirements set out by regulatory frameworks. However, their applicability is largely limited to structured inputs such as tabular data, and are not directly applicable to unstructured data such as natural language or images.
Recent advancements in large language models (LLMs) have substantially enhanced the capacity for semantic understanding and reasoning within natural language processing (Brown et al., 2020; Achiam et al., 2023). Building on these models, agentic systems (Shinn et al., 2023; Yao et al., 2023; Schick et al., 2023) have been proposed to address complex tasks involving reasoning over diverse types of data, and are seen as a very promising direction for future AI systems (Chen et al., 2023; Wang et al., 2024).
@@
Yet, despite their impressive capabilities, LLMs still suffer from several drawbacks—such as hallucinations, inconsistencies, unreliability, and difficulty to audit (Kryściński et al., 2019; Ji et al., 2023a; Tamkin et al., 2021). These issues constitute major obstacles to building reliable AI systems, limiting their adoption in sensitive or regulated domains. A growing body of work seeks to address these weaknesses by guiding LLM behavior toward more structured or self-consistent reasoning. Chain-of-Thought prompting Wei et al. (2022), for example, encourages multi-step reasoning by inserting explicit intermediate steps into prompts, improving consistency and sometimes reducing hallucinations. Other techniques—such as self-reflection Shinn et al. (2023); Madaan et al. (2023) and prompt optimization Zhou et al. (2022); Ji et al. (2023b); Yuksekgonul et al. (2024); Renze & Guven (2024)—introduce agentic feedback loops to iteratively refine model outputs. While these methods improve reliability in practice, they still rely on free-form text generation and implicit reasoning patterns that remain difficult to verify, audit, or formally constrain.
In this paper, we introduce a new type of classifier designed to combine the semantic reasoning capabilities of large language models (LLMs) with the transparency of decision trees. The idea is to follow a divide-and-conquer paradigm, allowing to reduce hallucination and reasoning errors by successive decomposition of the problem following a hierarchical decision logic.
@@
We therefore propose to structure LLM and agentic reasoning to address complex classification tasks through verifiable decision patterns on unstructured data such as images or texts, in the form of an agentic decision tree. Adapting the traditional Decision tree algorithms such as CART Breiman et al. (1984) and C4.5 Salzberg (1994), we propose ACT, an agentic decision tree where each node is defined as a natural language binary question over the input data, iteratively partitioning the data into two subsets at each step. The best splits are found using TextGrad Yuksekgonul et al. (2024), a prompt-refinement technique providing textual feedback to an agent, to minimize the Gini criterion Breiman et al. (1984) at each split. Ultimately, ACT takes the form of a decision tree (cf. Figure 1), each leaf leading to a class prediction after successive splitting.
The benefits of ACT are therefore the following:
@@
  • Transparent and structured decision process: ACT enforces a systematic, rule-based structure that aligns the model’s internal reasoning with its observable decision process. Its tree-based architecture provides fully traceable and interpretable decision paths, facilitating auditability, human oversight, and intervention—akin to the properties of traditional expert systems.
@@
  • Optimized decision process: we show that decomposing some tasks into subquestions allows to improve performances of LLMs without retraining. By automatically finding the most relevant questions at each node, ACT is able to surpass existing methods.
@@ 2 Problem Setup and Related Work
We consider a binary classification problem over an unstructured input space $\mathcal{X}$ (e.g., text or images), with labels $y\in\{0,1\}$. Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ sampled i.i.d. from an unknown distribution, the goal is to learn a classifier $f:\mathcal{X}\rightarrow\{0,1\}$ that is both accurate and interpretable. There is an abundant literature in XAI that focuses on defining desiderata for interpretable models Doshi-Velez & Kim (2017); Lipton (2018). Building on the regulatory requirements discussed in Section 1, we therefore aim to build a decision model satisfying the following properties:
@@
  • Contestability: the ability to contest an algorithmic-based decision by disputing the conditions it relies on Wachter et al. (2017); Venkatasubramanian & Alfano (2020); Lyons et al. (2021).
@@
Translating these desiderata into technical requirements is made especially difficult given the opaqueness and complexity of modern architectures such as LLMs and VLMs. There is a clear lack of literature support for what explanations satisfying these properties should look like in these contexts. As a result, in this work, we propose to define interpretability through the notion of explicit decision paths—structured sequences of semantically meaningful reasoning steps that can be inspected, understood, and verified by human users. Below, we ground this proposition in the existing XAI and LLM literatures.
Interpretable models and post-hoc explanations
@@
Interpretable models such as decision trees (Breiman et al., 1984; Salzberg, 1994) are traditionally viewed as a natural solution to meet the three aforementioned criteria. Offering transparent, rule-based decision paths that are easily understood and audited, their hierarchical structure allows users to trace predictions through explicit, human-readable splits, making them well-suited for regulated or high-stakes domains (Rudin, 2019). However, they are fundamentally limited to structured inputs, relying on predefined features and struggling to handle unstructured data such as raw text or images. Efforts to address interpretability in unstructured data contexts include post-hoc explanation techniques such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg & Lee, 2017) proposing explanations in the form of feature attributions at the pixel (for images) or word (for text) level. Although very popular, these methods have been criticized for often failing to capture the model’s true reasoning (Rudin, 2019; Laugel et al., 2019), consequently offering limited guarantees for auditability, consistency, or regulatory compliance.
More recent efforts in neural NLP have attempted to incorporate explanations directly into the model’s output, for instance through self-explanation (Nye et al., 2021) or agentic feedback loops (Madaan et al., 2023). However, these approaches rely on free-form natural language generation, which—although interpretable to humans—remains unverifiable and unstructured. As a result, they fall short of providing a consistent and auditable decision procedure.
Workflow and prompt optimization
@@
Research on building trustworthy LLM and VLM-based systems has largely focused on workflow design and prompt refinement. Approaches such as AutoPrompt (Shin et al., 2020), TextGrad (Yuksekgonul et al., 2024), and DSPy (Khattab et al., 2023) automatically optimize the prompt instructions to better align the LLM behavior with downstream tasks. These methods show that LLMs can be guided toward more accurate or systematic outputs without retraining, but do not provide explanation patterns. Techniques such as Chain-of-Thought prompting (Wei et al., 2022), ReAct (Yao et al., 2023), and Reflexion (Shinn et al., 2023) structure reasoning through intermediate steps or self-feedback, providing both performance increases and transparent decision patterns. However, their reliance on free-form text generation makes them still vulnerable to hallucinations, questioning their efficacy in auditing and verification contexts.
@@
We propose to address these challenges by enforcing a hierarchical tree-structured decision process over unstructured inputs. Defining each node as a natural-language question optimized through LLM feedback ensures both semantic adaptability and transparent reasoning paths. Our proposed framework, called Agentic Classification Tree (ACT), thus aims to combine the interpretability of decision trees with the semantic reasoning of LLMs, providing an accountable alternative to existing black-box prompt optimization methods.
@@ 3 Methodology
Table 1: Side-by-side comparison between traditional decision tree algorithms (CART, C4.5) and our proposed Agentic Classification Tree (ACT).
CART / C4.5 Agentic Classification Tree
Data type Tabular Unstructured
Node definition Numerical: binary threshold (both).
Categorical: CART binary subset; C4.5 multiway split.
Binary natural-language question queried to an LLM over the inputs.
Best split search Greedy exhaustive search over candidate splits. Iterative refinement of the question through prompt optimization (TextGrad) with LLM feedback.
Split criterion Gini (CART) or Information Gain Ratio (C4.5). (Gini or IG) combined with semantic purity analysis.
Inference method Each input follows the tree based on feature values. Each input is queried by the LLM at each node and routed according to its answers.
To build a structured agentic workflow in the form of a classification tree, we take inspiration from traditional decision tree algorithms such as CART (Breiman et al., 1984) and C4.5 (Salzberg, 1994). Table 1 provides a side-by-side comparison between traditional decision trees and our proposed ACT method.
@@
Classical decision-tree algorithms such as CART (Breiman et al., 1984) and C4.5 (Salzberg, 1994) recursively partition a dataset based on feature values. At each node, a split is defined by either a threshold on a numerical feature or a subset of categorical values. The quality of each split is measured by an impurity criterion—typically Gini impurity (CART) or Information Gain ratio (C4.5)—and the algorithm greedily selects the split that yields the greatest impurity reduction, continuing recursively until a stopping criterion is met (e.g., maximum depth or node purity).
These methods are efficient and highly interpretable for structured, tabular data, but cannot operate directly on unstructured inputs such as text or images. This limitation motivates our proposed ACT, which replaces feature-based splits with optimized natural language queries evaluated by an LLM.
3.2 Proposed Approach: Agentic Classification Tree (ACT)
The Agentic Classification Tree (ACT) is designed to extend classical decision tree methods to unstructured data. Like CART and C4.5, ACT recursively partitions the input space, but instead of relying on numeric or categorical features, it leverages large language models (LLMs) to define natural-language queries that guide data splits.
Node definition
Each node $x$ in ACT is defined by a natural-language prompt $p$ to semantically partition data via an LLM. Given an instance $x_{i}$, the LLM is queried with $p$ and responds with a binary answer (yes or no), determining the branch assignment. This process defines the following split:
$\mathcal{D}_{L}^{x}=\{(x_{i},y_{i})\in\mathcal{D}^{x}\mid f_{\text{split}}(p,x_{i})=\texttt{yes}\},\quad\mathcal{D}_{R}^{x}=\mathcal{D}^{x}\setminus\mathcal{D}_{L}^{x}.$ (1)
where $f_{\text{split}}(p,x_{i})$ denotes the LLM’s response to prompt $p$ on input instance $x_{i}$. The resulting subsets $\mathcal{D}_{L}^{x}$ and $\mathcal{D}_{R}^{x}$ thus reflect the semantic partition induced by the model at node $x$.
To initialize the splitting process, we begin with a neutral, generic question such as:
"Based on the provided example, does it belong to the positive class? (yes/no)"
This provides an unbiased starting point. Since $f_{\text{split}}$ is neither trained on the task nor informed of the target labels, the initial partition is not expected to be meaningful. This is by design: withholding task-specific information ensures that decision prompts are derived from the training data rather than from prior knowledge. In contrast to approaches such as TextGrad (Yuksekgonul et al., 2024), which typically begin with a task-specific question and iteratively refine it, ACT starts from a neutral prompt and discovers semantically meaningful questions from scratch through data-driven refinement.
Best split criterion

### v3

@@ Abstract
When used in high-stakes settings, AI systems are expected to produce decisions that are transparent, interpretable and auditable—a requirement increasingly expected by regulations. Decision trees such as CART provide clear and verifiable rules, but they are restricted to structured tabular data and cannot operate directly on unstructured inputs such as text. In practice, large language models (LLMs) are widely used for such data, yet prompting strategies such as chain-of-thought or prompt optimization still rely on free-form reasoning, limiting their ability to ensure trustworthy behaviors. We present the Agentic Classification Tree (ACT), which extends decision-tree methodology to unstructured inputs by formulating each split as a natural-language question, refined through impurity-based evaluation and LLM feedback via TextGrad. Experiments on text benchmarks show that ACT matches or surpasses prompting-based baselines while producing transparent and interpretable decision paths. Code: https://github.com/axa-rev-research/ACT
Machine Learning, ICML
@@ 1 Introduction
Unstructured Data: Tuberculosis diagnosis based on symptoms descriptions Patient 1: ”My main symptom is a runny nose with increased sweating, pain, fever, skin rashes, nasal congestion, sore throat, muscle pain, loss of appetite, and chills. The heavy pain is in my left temple … ” Patient 2: ”I am coughing blood with pain (heavy in my left temple), skin lesions, nasal congestion, extreme fatigue affecting activities, diffuse muscle pain, loss of appetite, chills, a heavy pain on my left temple … ” Refer to caption Refer to caption
{forest}
Figure 1: Example ACT for tuberculosis (TB) diagnosis using unstructured, free-text patient descriptions. A decision tree is automatically learned, with each node containing a binary natural-language question, autonomously discovered via recursive prompt refinement to maximize label separation at each split. At inference, given an input sample, these questions are answered by an LLM from the root node to the leaves. The final classification (TB or Not TB) corresponds to the majority label of training examples described by each leaf.
@@
As AI systems become increasingly integrated into high-stakes domains such as healthcare, education, legal decision-making and finance, the need for transparency, interpretability and auditability in AI decision-making has intensified. These requirements are grounded not only in practical considerations, but also in legal obligations: recent governance frameworks and regulations such as the EU AI Act the NIST AI Risk Management Frameworkemphasize that AI-based decisions in high-stakes scenarios must be explainable and subject to human oversight. In this context, there is a growing need for models whose decision processes can be inspected, verified and understood not only by technical experts but also by stakeholders, auditors and regulators.
@@
Historically, complex decision-making tasks in high-stakes domains have been addressed by interpretable AI systems such as expert systems (jackson1986introduction; shortliffe1986medical; leonard1993detecting; talebzadeh1995countrywide) and decision trees (DBLP:books/wa/BreimanFOS84). These models allow users to trace each decision through a sequence of explicit, human-understandable rules (slack2019assessing), directly supporting the requirements set out by regulatory frameworks [^1] https://airc.nist.gov/airmf-resources/playbook/. However, their applicability is largely limited to structured inputs such as tabular data and are not directly applicable to unstructured data such as natural language or images.
@@
Recent advancements in large language models (LLMs) have substantially enhanced the capacity for semantic understanding and reasoning within NLP  (brown2020language; achiam2023gpt). Building on these models, agentic systems (shinn2023reflexion; yao2023react; schick2023toolformer) have been proposed to address complex tasks involving reasoning over diverse types of data, and are seen as a very promising direction for future AI systems (chen2023agentverse; wang2024survey). Yet, despite their impressive capabilities, LLMs still suffer from several drawbacks—such as hallucinations, unreliability and difficulty to audit (kryscinski2019evaluating; ji2023survey; tamkin2021understanding). These issues constitute major obstacles to building reliable AI systems, limiting their adoption in sensitive or regulated domains. A growing body of work seeks to address these weaknesses by guiding LLM behavior toward more structured or self-consistent reasoning. Chain-of-Thought prompting (wei2022chain), for example, encourages multi-step reasoning, improving consistency and sometimes reducing hallucinations. Other techniques, such as self-reflection (shinn2023reflexion; madaan2023self) and prompt optimization (zhou2022large; ji2023towards; yuksekgonul2024textgrad; renze2024self), introduce agentic feedback loops to iteratively refine model outputs. While these methods improve reliability, they still rely on free-form text generation and implicit reasoning that remain difficult to verify or audit.
In this paper, we introduce a new type of classifier designed to combine the semantic reasoning capabilities of LLMs with the transparency of decision trees. The idea is to follow a divide-and-conquer paradigm by decomposing a difficult classification question (e.g., Does the patient have disease X?) into a series of easier questions (e.g., Is symptom Y present?), as shown in Figure 1. We therefore propose to structure LLM and agentic reasoning as an agentic decision tree, addressing complex classification tasks on unstructured data through verifiable decision patterns while reducing reasoning errors via hierarchical problem decomposition.
@@
Adapting traditional decision tree algorithms such as CART (DBLP:books/wa/BreimanFOS84) and C4.5 (quinlan1993c4), we propose the Agentic Classification Tree (ACT). In ACT, each node is defined as a binary natural-language question over the input data, iteratively partitioning the data into two subsets at each step. The best splits are found using TextGrad (yuksekgonul2024textgrad), a prompt-refinement technique providing textual feedback to an agent, to minimize the Gini criterion (DBLP:books/wa/BreimanFOS84) at each split. Ultimately, ACT takes the form of a decision tree (cf. Figure 1), each leaf leading to a class prediction after successive splitting. The benefits of ACT are therefore the following:
@@
  • Transparent and structured decision process: ACT enforces a systematic, rule-based structure that aligns the model’s internal reasoning with its observable decision process. Its tree-based architecture provides fully traceable and interpretable decision paths, facilitating auditability, human oversight and intervention.
@@
  • Optimized decision process: we show that decomposing some tasks into subquestions improves the performance of LLMs without retraining.
@@ 2 Problem Setup and Related Work
We consider a binary classification problem over an unstructured input space $\mathcal{X}$ (e.g., text or images), with labels $y\in\{0,1\}$. Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ sampled i.i.d. from an unknown distribution, the goal is to learn a classifier $f:\mathcal{X}\rightarrow\{0,1\}$ that is both accurate and interpretable. There is an abundant literature in XAI that focuses on defining desiderata for interpretable models (doshi2017towards; lipton2018mythos). Building on the regulatory requirements discussed in Section 1, we therefore aim to build a decision model satisfying the following properties:
@@
  • Contestability: the ability to contest an algorithm-based decision by disputing the conditions it relies on (wachter2017counterfactual; venkatasubramanian2020philosophical; lyons2021conceptualising).
@@
Translating these desiderata into technical requirements is challenging given the opacity of modern architectures such as LLMs and VLMs, with limited literature guidance on suitable explanations in these contexts. As a result, in this work, we propose to define interpretability through the notion of explicit decision paths—structured sequences of semantically meaningful reasoning steps that can be inspected, understood and verified by human users. Below, we ground this proposition in the existing XAI and LLM literatures.
Interpretable models and post-hoc explanations.
@@
Interpretable models such as decision trees (DBLP:books/wa/BreimanFOS84; quinlan1993c4) are traditionally viewed as a natural solution to meet the three aforementioned criteria. Their transparent, rule-based structure allows users to trace predictions through human-readable splits, making them well-suited for regulated or high-stakes domains (rudin2019stop). However, they are fundamentally limited to structured inputs, relying on predefined features and struggling to handle unstructured data such as raw text or images. Efforts to address interpretability in unstructured data contexts include post-hoc explanation techniques such as LIME (ribeiro2016should) and SHAP (lundberg2017unified) proposing explanations in the form of feature attributions at the pixel (for images) or the word level (for text). Although very popular, these methods have been criticized for often failing to capture the model’s true reasoning (rudin2019stop; laugel2019dangers), consequently offering limited guarantees for auditability, consistency or regulatory compliance. More recent efforts in neural NLP have attempted to incorporate explanations directly into the model’s output, for instance through self-explanation (nye2021show) or agentic feedback loops (madaan2023self). However, these approaches rely on free-form natural language generation, which—although interpretable to humans—remains unverifiable and unstructured. As a result, they fall short of providing a consistent and auditable decision procedure.
Workflow and prompt optimization.
@@
Research on building trustworthy LLM and VLM-based systems has largely focused on workflow design and prompt refinement, with approaches such as AutoPrompt (shin2020autoprompt), TextGrad (yuksekgonul2024textgrad) and DSPy (khattab2023dspy) automatically optimizing prompt instructions to align LLM behavior with downstream tasks. These methods show that LLMs can be guided toward more accurate or systematic outputs without retraining, but do not provide explanation patterns. Techniques such as Chain-of-Thought prompting (wei2022chain), ReAct (yao2023react), and Reflexion (shinn2023reflexion) structure reasoning through intermediate steps or self-feedback, providing both performance increases and transparent decision patterns. However, their reliance on free-form text generation makes them still vulnerable to hallucinations, questioning their efficacy in auditing and verification contexts.
@@
We propose to address these challenges by enforcing a hierarchical tree-structured decision process over unstructured inputs. Defining each node as a natural-language question optimized through LLM feedback ensures both semantic adaptability and transparent reasoning paths. Our proposed framework, called Agentic Classification Tree, thus aims to combine the interpretability of decision trees with the semantic reasoning of LLMs, providing an accountable alternative to existing black-box prompt optimization methods.
@@ 3 Methodology
To build a structured agentic workflow in the form of a classification tree, we take inspiration from traditional decision tree algorithms such as CART (DBLP:books/wa/BreimanFOS84) and C4.5 (quinlan1993c4). Table 1 provides a side-by-side comparison between them and our proposed ACT method.
@@
Table 1: Comparison between traditional decision trees and ACT.
CART / C4.5 ACT (ours)
Data type Tabular features Unstructured inputs (text, images, etc.)
Node definition Numerical: binary threshold, Categorical: binary subset or multiway split Binary natural- language question queried by an LLM over the inputs.
Best split search Greedy search over candidate feature splits. Iterative question refinement via TextGrad.
Split criterion Gini (CART) or Information Gain Ratio (C4.5). Gini impurity combined with semantic signals.
Inference Each input follows the tree based on feature values. Each input is queried by the LLM at each node and routed according to its (yes/no) answers.
Classical decision-tree algorithms recursively partition a dataset based on feature values. At each node, a split is defined by either a threshold on a numerical feature or a subset of categorical values. The quality of each split is measured by an impurity criterion—typically Gini impurity (CART) or Information Gain ratio (C4.5)—and the algorithm greedily selects the split that yields the greatest impurity reduction, continuing recursively until a stopping criterion is met. These methods are efficient and interpretable  (molnar2020interpretable) for structured, tabular data, but cannot operate directly on unstructured inputs such as text or images. This limitation motivates our proposed ACT, which replaces feature-based splits with optimized natural language queries.
3.2 Proposed Approach: Agentic Classification Tree
ACT is designed to extend classical decision tree methods to unstructured data by replacing numeric or categorical feature splits with LLM-generated natural-language questions that recursively partition the input space.
Node definition.
Each node $n$ in ACT is defined by a prompt $p$ to semantically partition data via an LLM. Given an instance $x_{i}$, the LLM is queried with $p$ and responds with a binary answer (yes or no), determining the branch assignment. This process defines the following split:
$\begin{gathered}\mathcal{D}_{L}^{n}=\{(x_{i},y_{i})\in\mathcal{D}^{n}\mid f_{\text{split}}(p,x_{i})=\texttt{yes}\},\\ \mathcal{D}_{R}^{n}=\mathcal{D}^{n}\setminus\mathcal{D}_{L}^{n}.\end{gathered}$
where $f_{\text{split}}(p,x_{i})$ denotes the LLM’s response to prompt $p$ on input $x_{i}$. The resulting subsets $\mathcal{D}_{L}^{n}$ and $\mathcal{D}_{R}^{n}$ thus reflect the semantic partition induced by the model at node $n$. The splitting process is initialized with a neutral, generic question and in our experiments we use:
Based on the provided example, does it belong to the positive class? (yes/no)
Since $f_{\text{split}}$ is neither trained on the task nor informed of the target labels, the initial partition is not expected to be meaningful. This is by design: withholding task-specific information ensures that decision prompts are derived from the training data rather than from prior knowledge. Unlike TextGrad (yuksekgonul2024textgrad), which begins with a task-specific question and iteratively refines it, ACT starts from a neutral prompt and discovers meaningful questions through data-driven refinement.
Best split criterion.